OpenStudy (anonymous):

Evalute 2n2 − 7n − 4 given n(n − 2) = 15. (smaller value) (larger value)

OpenStudy (anonymous):

Evaluate \[2n^2 - 7n - 4\] given \[n^2 - 2n -15 = 0\]

OpenStudy (anonymous):

n^2 - 2n - 15 = (n-5) (n +3) = 0 n = 5, -3

OpenStudy (anonymous):

substite the values of n to the evaluate equation

OpenStudy (anonymous):

n^2 -2n = 15 n^2-2n-15=0 used factorisation (n-5)(n+3)=0 n=5, n=-3 and then substitute into the other equation both answers.

OpenStudy (anonymous):

i did this one earlier on answers 11 and 35 if my memory serves me right

OpenStudy (anonymous):

Yes, 11 and 35 are the right answers after all those steps have been done.